# The Impact of Auxiliary Patient Data on Automated Chest X-Ray Report Generation and How to Incorporate It

### v1

@@ 4.2 Tabular Data
An example table and its conversion to embeddings is shown in Figure 3. To convert an exam’s tabular data to embeddings, columns were designated as value, category, text, or time columns. Value columns contained numeric data, while category columns contained categorical data. Datum from value and category columns were grouped by their time delta, with each group forming a feature vector. The feature vector initially consisted of zeros. Values and categories from the group were then used to set its values based on indices determined by a lookup table. For value columns, the lookup table determined the index where the numeric value was placed. For category columns, it determined which indices were activated (set to 1).
@@
Next, the feature vector was passed through an FNN $f(\bm{X}_{i}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$ to form the embedding, where $\bm{X}_{i}\in\mathbb{R}^{|U_{C}|,|L_{i}|}$ are the grouped features, $\bm{W}_{\text{1}}\in\mathbb{R}^{|L_{i}|,2048}$ and $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $L_{i}$ is a lookup table, and $i$ designates the table. Each table has a unique FNN and lookup table. Rows for a value column always had a unique time, preventing multiple values from the same column in a group. We investigated alternatives to form the value-category embeddings in Section 5. The described framework was found to be the most efficient. Columns with a high cardinality were set as text columns. Text embeddings were formed via the decoder’s tokenizer and token embeddings. Text embeddings were given the time delta embedding from their respective row. The column designation for each table in Figure 1 is described in the Appendix B.
4.3 Report Section Embeddings
@@
Here, we consider five sections of the radiology report: the findings, impression, indication, history, and comparison sections. The findings and impression sections serve as the ground truth to be generated. The remainder form part of the patient data embeddings. The indication section explains the reason for the exam, such as symptoms or suspected conditions. The history section provides relevant medical history, such as past conditions and treatments. The comparison section mentions any prior exams used to identify changes over time. These sections provide context that guides the interpretation of the exam, influencing the content of the findings and impression sections. The embeddings were formed via the decoder’s tokenizer and token embeddings. Of these, the history and comparison sections have not been investigated for CXR report generation. The comparison section was used only when prior exams were considered.
@@
Images 26.00 29.24 58.87 24.10 12.24 14.35 24.34 6.33 272.4
Patient Emergency Department (ED) data (MIMIC-IV-ED)
Images + ED stays 26.10 29.47 60.65 24.17 12.39 14.52 24.50 6.36 273.4
Images + triage 26.46 31.27 63.06 24.29 12.32 14.66 24.58 6.44 278.9
Images + vital signs 26.47 31.72 63.39 24.32 13.16 14.61 24.74 6.47 274.7
Images + medicine reconciliation 26.86 31.37 63.98 24.52 12.77 14.90 24.85 6.60 343.5
Images + medicine administration 26.15 29.47 59.21 24.25 12.30 14.44 24.47 6.38 273.0
Patient additional radiology data (MIMIC-CXR)
Images + indication 26.94 32.13 65.43 24.74 14.16 15.19 25.16 7.02 279.5
Images + history 27.00 31.88 65.06 25.05 14.32 15.30 25.48 7.33 277.0
Images + metadata 26.34 29.63 59.55 24.37 12.40 14.55 24.50 6.43 273.4
@@
Images + $h=1$ 26.98 31.42 63.98 24.65 12.65 15.11 25.03 6.78 558.9
Images + $h=1$ + comparison 26.76 31.55 64.20 24.42 13.36 15.03 24.82 6.74 563.4
Images + $h=2$ 26.67 30.48 61.27 24.53 13.60 14.94 24.85 6.72 810.6
Images + $h=2$ + comparison 26.20 30.19 61.24 24.05 12.43 14.80 24.55 6.58 815.0
Images + $h=3$ 26.47 29.96 59.95 24.14 12.90 14.94 24.66 6.65 1037.1
Images + $h=3$ + comparison 26.14 30.09 60.51 23.90 13.22 14.87 24.56 6.64 1041.5
All effective sources (no medicine administration, metadata, or comparison)
Images + effective sources ($h=0$) 27.11 32.23 64.80 25.07 14.48 15.15 25.40 7.07 365.0
Images + effective sources ($h=1$) 26.78 31.83 63.85 24.75 14.10 15.15 25.25 7.01 651.7
@@
    - medicine reconciliation 26.78 32.81 65.60 24.84 14.44 15.21 25.33 7.19 293.9
    - ED stays 26.94 31.56 64.87 25.02 14.08 15.14 25.37 7.09 364.0
    - triage 27.15 32.45 65.18 25.15 14.80 15.27 25.54 7.25 358.5
    - vital signs 27.27 31.78 65.44 25.14 14.07 15.35 25.49 7.22 362.6
    - indication 26.89 31.25 64.65 24.99 13.87 15.07 25.39 7.00 357.9
    - history 26.96 31.87 64.02 24.86 14.60 15.10 25.24 7.04 360.3
    - time delta 27.17 32.11 65.10 25.18 14.64 15.24 25.54 7.16 365.0
4.4 Experiment Setup
Our multimodal language model, illustrated in Figure 2, is based on CXRMate-RRG24; it features a Llama decoder and the UniFormer as the image encoder. The training procedure for our model involved three stages: (1) initial training on the MIMIC-CXR training set using only images as input with Teacher Forcing (TF) (Williams and Zipser, 1989), (2) further training on the dataset described in Section 1 with the inputs detailed in Table 1, again using TF, and (3) reinforcement learning on the same dataset through self-critical sequence training (SCST) Rennie et al. (2017) (only for Table 2). Our evaluation metrics included three that capture the semantics of radiology reporting — RadGraph-F1 (RG), CheXbert-F1 (CX), and CXR-BERT (CB) — as well as five natural language generation metrics: BERTScore-F1 (BS), CIDEr (C), METEOR (M), ROUGE-L (R-L), and BLEU-4 (B4). Comprehensive details on the model architecture, training procedure, significance testing, and comparison methods are provided in Appendix C.
@@ 4.4 Experiment Setup
Model Train samples RG CX CB BS C M R-L B4
EMNLI Miura et al. (2021) 152 173 exams 32.8 28.9 66.6 24.4 19.4 17.1 28.1 8.9
CMN Chen et al. (2021) 270 790 images 25.3 24.3 49.4 19.7 16.9 15.1 26.4 7.6
TranSQ Kong et al. (2022) 368 960 images 29.8 30.4 62.3 20.4 14.9 17.6 22.6 7.9
RGRG Tanida et al. (2023) 166 512 images 23.2 22.8 37.9 23.4 7.6 12.4 21.1 5.4
CvT2DistilGPT2 Nicolson et al. (2023) 270 790 images 25.8 29.3 59.8 24.8 20.9 16.0 27.3 8.8
RaDialog Pellegrini et al. (2023) 276 778 images 26.8 38.4 60.7 26.2 14.6 14.7 25.4 6.9
MedXChat Yang et al. (2023) 270 790 images 22.6 13.1 21.3 19.3 9.8 14.3 23.2 7.0
CXR-LLaVA-v2 Lee et al. (2024) 193 513 images 20.7 20.7 44.1 23.6 5.2 11.3 19.9 2.7
CXRMate Nicolson et al. (2024a) 125 395 exams 28.8 33.9 71.3 30.5 22.4 17.7 28.1 9.7
CXRMate-RRG24 Nicolson et al. (2024b) 550 395 exams 30.4 31.2 58.2 31.0 20.6 16.7 27.5 9.1
Ours 76,398 exams 33.7 35.1 79.1 35.8 24.1 19.1 30.6 11.9
Figure 4: Case study demonstrating how incorporating a diverse set of patient data can aid with report generation.
@@ 5 Results & Discussion
The impact of different patient data sources on the performance of CXR report generation is summarised in Table 1. This analysis identifies which additional data sources enhance performance compared to using only images.
Significant improvements were observed by incorporating either the ED stays, triage, medicine reconciliation, or vital signs data from MIMIC-IV-ED dataset. Notably, the ED data markedly improved scores on the radiology report metrics (RG, CX, and CB). The medicine administration table did not significantly improve the scores overall, likely due to its infrequent occurrence in the exams (37%). (However, as shown in Table 4, it significantly improves performance when evaluated solely on exams that include a medicine administration table.) These findings demonstrate that ED patient data can enhance the diagnostic accuracy of CXR report generation.
@@
[…]
@@
Next, we further trained ‘Images + effective sources (h = 0) - medicine reconciliation’ with reinforcement learning, as described in Subsection 4.4. This model, denoted as ‘Ours’ in Table 2, was compared to other benchmark CXR report generation models in the literature that included MIMIC-CXR in their training data. Despite having substantially fewer training samples than the other models, our model significantly outperformed them on CXR-BERT, BERTScore-F1, METEOR, ROUGE-L, and BLEU-4. This demonstrates the impact of incorporating a more comprehensive set of patient data on CXR report generation.
A case study is presented in Figure 4 demonstrating how a diverse set of patient data can impact report generation. Here, the first model is given the image only, and fails to identify key findings that the radiologist noted in their report. The second model is given the additional patient data available for this exam; the indication section and triage data. Hypoxia, as indicated by the low oxygen saturation (‘o2sat’), along with the elevated respiratory rate (‘resprate’) and systolic blood pressure (‘SBP’), are consistent with the physiological responses to pulmonary edema. Given this, the second model was able to identify the moderate pulmonary edema, echoing the radiologist’s findings.
@@
Table 3 compares different methods for converting value and category columns into embeddings. This evaluation includes images, the triage table, and the medicine reconciliation table, as these tables contain multiple value and category columns. The aforementioned method of producing embeddings by grouping data from value and category columns (‘Grouped embeddings’) is compared to two other methods. The first is separate embeddings for each datum, where each value column datum is separately transformed using the previously described FNN, while each category column datum is converted to an embedding using a learnable weight matrix, akin to how token embeddings are produced (‘Separate embeddings’). The second method modifies ‘Separate embeddings’ by instead converting the value column data to text and using the decoder’s tokenizer and token embeddings (‘Values-to-text, categories-to-tokens’). The results indicate that the grouped embeddings method generally works best and is useful for encoding heterogeneous patient data for multimodal models.
@@ 6 Conclusion
This paper demonstrates the value of incorporating diverse patient data into automated CXR report generation. By integrating patient data from the MIMIC-CXR and MIMIC-IV-ED datasets, we have shown significant improvements in the diagnostic accuracy of generated radiology reports. Our empirical evaluation uncovers new sources of patient information that enhance CXR report generation, including data from ED stays, triaging information, aperiodic vital signs, medications, and the history section of radiology reports. We present specific methods to convert multimodal patient data into embeddings for a language model, encompassing numerical, categorical, textual, temporal, and image data. We encourage further research and experimentation using our released dataset splits, code, and model checkpoints to explore innovative methods for multimodal patient data integration, with the ultimate goal of enhancing diagnostic accuracy and patient care.
[…]
Embeddings CX RG CB BS
@@
Baseline 25.81 29.00 59.04 23.85
Images + triage + medicine reconciliation
Grouped embeddings 26.72 31.69 64.01 24.38
Separate embeddings 25.32 25.28 46.29 23.51
Values-to-text, categories-to-embeddings 26.46 30.70 58.62 24.58

### v2

@@ 4.2 Patient Data Embeddings: Tabular Data
An example table and its conversion to embeddings is shown in Figure 3. The columns of each table were designated as value, category, text, or time columns. Value columns contained numeric data, while category columns contained categorical data. To convert an exam’s tabular data to embeddings, data from value and category columns were grouped by their time delta, where each group formed a feature vector. The feature vector initially consisted of zeros. Values and categories from the group were then used to set its values based on indices determined by a lookup table. For value columns, the lookup table determined the index where the numeric value was placed. For category columns, it determined which indices were activated (set to 1).
@@
Next, the feature vector was passed through an FNN $f(\bm{X}_{i}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$ to form the embedding, where $\bm{X}_{i}\in\mathbb{R}^{|U_{C}|,|L_{i}|}$ are the grouped features, $U_{C}$ is the set of unique time deltas, $\bm{W}_{\text{1}}\in\mathbb{R}^{|L_{i}|,2048}$ and $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $L_{i}$ is a lookup table, and $i$ designates the table. Each table has a unique FNN and lookup table. Rows for a value column always had a unique time, preventing multiple values from the same column in a group. We investigated alternatives to form the value-category embeddings in Section 6. The described framework was found to be the most efficient. Columns with a high cardinality were set as text columns. Text embeddings were formed via the decoder’s tokenizer and token embeddings. Text embeddings were given the time delta embedding from their respective row. The column designation for each table in Figure 1 is described in the Appendix C.
4.3 Patient Data Embeddings: Report Sections
@@
Here, we consider five sections of the radiology report: the findings, impression, indication, history, and comparison sections. The findings and impression sections serve as the ground truth to be generated. The remainder form part of the patient data embeddings. The indication section explains the reason for the exam, such as symptoms or suspected conditions. The history section provides relevant medical history, such as past conditions and treatments. The comparison section mentions any prior exams, which are used to capture disease progression. These sections provide context that guides the interpretation of the exam, influencing the content of the findings and impression sections. The embeddings were formed via the decoder’s tokenizer and token embeddings. Of these, the history and comparison sections have not been investigated for CXR report generation. The comparison section was used only when prior exams were considered.
@@
Images 24.54 30.10 59.25 35.16 24.26 25.91 4.75 272.4
Patient emergency department data (MIMIC-IV-ED)
Images + ED stays 24.20 29.55 58.37 34.64 24.06 25.77 4.66 273.4
Images + triage 24.59 31.33 62.79 35.78 24.40 25.96 4.76 278.9
Images + vital signs 24.23 30.61 60.61 35.15 24.04 25.86 4.70 274.7
Images + reconciled medicines 25.10 32.05 64.70 36.32 24.71 26.29 4.93 355.6
Images + administered medicines 24.22 30.40 60.13 34.85 23.97 25.61 4.58 273.0
Patient radiology data (MIMIC-CXR)
Images + indication 25.01 32.78 65.49 35.88 24.73 26.32 5.15 279.5
Images + history 24.88 31.66 63.91 35.76 24.91 26.70 5.54 277.0
Images + metadata 24.07 30.42 59.75 34.79 23.86 25.59 4.58 273.4
@@
Images + $h=1$ 24.71 30.98 62.60 35.81 24.38 26.00 4.82 603.0
Images + $h=2$ 24.56 31.43 62.09 35.43 24.04 25.80 4.84 878.1
Images + $h=3$ 24.50 30.73 59.89 35.21 24.03 25.82 4.70 1134.3
Images + $h=1$ + comparison 24.92 31.46 62.93 35.84 24.34 26.03 4.89 607.4
Images + $h=2$ + comparison 24.52 31.01 61.36 34.89 23.90 25.62 4.72 882.6
Images + $h=3$ + comparison 24.31 30.93 60.10 34.35 23.31 25.39 4.72 1138.8
All effective sources (triage, reconciled medicines, indication, and history)
Images + effective sources ($h=0$) 25.52 32.49 65.93 36.26 25.16 26.81 5.34 373.9
Images + effective sources ($h=1$) 25.11 31.14 61.19 35.80 24.95 26.45 5.21 704.5
Images + effective sources ($h=1$ + comparison) 25.05 30.68 60.99 35.94 24.94 26.48 5.24 709.0
@@
    - triage 25.65 32.85 65.38 36.33 25.25 26.75 5.33 367.4
    - reconciled medicines 25.43 32.48 65.63 36.42 25.23 26.86 5.40 290.7
    - indication 25.46 32.92 65.69 36.41 25.21 26.79 5.36 366.7
    - history 25.41 32.53 65.82 36.65 25.12 26.72 5.37 369.2
    - time delta 25.31 33.03 65.72 36.17 25.10 26.75 5.34 373.9
5 Experiment Setup
Our multimodal language model, illustrated in Figure 2, is based on CXRMate-RRG24; it features a Llama decoder and the UniFormer as the image encoder. The training procedure for our model involved three stages: (1) initial training on the MIMIC-CXR training set using only images as input with Teacher Forcing (TF) (Williams and Zipser, 1989), (2) further training on the dataset described in Section 1 with the inputs detailed in Table 1, again using TF, and (3) reinforcement learning on the same dataset through self-critical sequence training (SCST) Rennie et al. (2017) (only for Table 2). Our evaluation metrics included four that capture the semantics of radiology reporting — RadGraph-F1 (RG), CheXbert-F1 (CX), CXR-BERT (CB), and GREEN (G) — as well as three natural language generation metrics: BERTScore-F1 (BS), ROUGE-L (R-L), and BLEU-4 (B4). We also propose a metric that measures n-gram repetition rate, namely the absence of repeated n-grams (ARN). Comprehensive details on ARN and the other metrics, the model architecture, training procedure, significance testing, and comparison methods are provided in Appendix D.
@@ 5 Experiment Setup
Model Train samples RG CX CB G BS R-L B4 ARN
EMNLI Miura et al. (2021) 152 173 exams 29.1 28.9 66.6 41.5 24.4 29.3 4.1 95.1
CMN Chen et al. (2021) 270 790 images 23.6 24.3 49.4 36.6 19.7 27.8 4.0 99.3
TranSQ Kong et al. (2022) 368 960 images 28.7 30.4 62.3 38.2 20.4 23.3 4.1 98.5
RGRG Tanida et al. (2023) 166 512 images 22.9 22.8 37.9 31.1 23.4 22.0 3.7 96.5
CvT2DistilGPT2 Nicolson et al. (2023) 270 790 images 23.9 29.3 59.8 37.0 24.8 28.6 5.4 99.0
RaDialog Pellegrini et al. (2023) 276 778 images 24.4 38.4 60.7 34.9 26.2 26.7 4.8 94.4
MedXChat Yang et al. (2023) 270 790 images 21.0 13.1 21.3 31.4 19.3 23.8 4.0 97.9
CXR-LLaVA-v2 Lee et al. (2024) 193 513 images 19.4 20.7 44.1 24.0 23.6 21.1 1.7 99.7
CXRMate Nicolson et al. (2024a) 125 395 exams 26.5 33.9 71.3 40.3 30.5 29.1 7.5 98.2
CXRMate-RRG24 Nicolson et al. (2024b) 550 395 exams 28.9 31.2 58.2 40.2 31.0 28.7 6.6 97.7
Images + effective sources ($h=0$) 76 398 exams 25.1 29.6 66.0 36.9 29.4 27.8 5.8 98.5
   + RL (CXR-BERT + BERTScore reward) 76 398 exams 30.4 35.7 79.1 41.6 37.2 31.6 8.7 93.5
    + reward per section 76 398 exams 30.1 33.7 78.3 41.6 37.5 32.2 8.4 94.6
     + ARN reward 76 398 exams 30.2 33.6 78.0 40.7 37.3 31.9 7.6 99.3
Figure 4: Case study demonstrating how incorporating auxiliary patient data can aid with report generation.
@@ 6 Results & Discussion
The impact of different patient data sources on the performance of CXR report generation is summarised in Table 1. This analysis identifies which additional data sources improve performance compared to using only images. Significant improvements were observed by incorporating either triage or reconciled medicine data from MIMIC-IV-ED dataset. Notably, this data markedly improved scores on the radiology report metrics (RG, CX, CB, and G). These findings demonstrate that ED patient data can improve the diagnostic accuracy of CXR report generation. Aperiodic vital sign and administered medicine data did not significantly improve the scores overall, likely due to their frequency of occurrence in the exams (62% and 37%, respectively). However, as shown in Table F.1, a significant improvement in performance was attained when evaluated solely on exams that include an aperiodic vital sign table.
@@
Next, we combined all effective sources of patient data (those providing a significant improvement). This included ‘triage’, ‘reconciled medicines’, ‘indication’, and ‘history’. The best performance was observed with no prior exams ($h=0$), indicating that using any prior exams in combination with other sources is detrimental with our model, possibly due to attention dilution. With $h=0$, the combination of all effective sources outperformed each individual source. We then conducted an ablation study using ‘Images + effective sources ($h=0$)’, which demonstrated that removing any individual patient data source did not result in a significant change in performance.
@@
Following this, we further trained ‘Images + effective sources (h = 0)’ with reinforcement learning (RL), as described in Subsection 5. Its performance is shown in Table 2; a CXR-BERT and BERTScore composite reward was used, which demonstrates a marked improvement for each metric, except ARN. The low ARN indicates that this reward introduced repetitions. We also propose to calculate the reward separately for the findings and impression section, as described in Appendix E. While this produces similar results for the findings section, as shown in Table 2, this significantly improves the scores on the impression section section, as shown in Table E.1. Finally, we incorporate ARN into the composite reward. This effectively reduces repetitions, as evidenced by the improved ARN, albeit with a slight trade-off in the other metrics. Compared to other benchmark CXR report generation models in the literature that included MIMIC-CXR in their training data, our model significantly outperformed them on multiple metrics in Table 2, despite having substantially fewer training samples. This demonstrates the impact of incorporating auxiliary patient data on CXR report generation.
Figure 4 demonstrates how auxiliary patient data enhances CXR report generation. Mild pulmonary edema was identified only when this data was incorporated. The patient’s low oxygen saturation, chief complaint of congestive heart failure (CHF) — a common cause of pulmonary edema — and reconciled medicines (Furosemide, Metoprolol Tartrate, Lisinopril, Spironolactone) indicate active management of fluid overload and cardiac dysfunction, all pointing to pulmonary edema. This supplementary evidence allowed the model to corroborate the imaging findings and identify pulmonary edema.
In Appendix G, we perform an error analysis to assess the influence of auxiliary patient data on the generated reports. Our findings show that incorporating auxiliary patient data increases the AUC for 10 out of the 14 CheXpert labels (Figure E.1), demonstrating its utility across multiple pathologies. Additionally, we analysed its impact on the generated reports for eight exams, with the following key observations:
True positives ($n=2$): The model utilised supportive auxiliary patient data effectively. (See Appendix G.2.1 and G.2.2.)
False positives ($n=2$): The model was misled by confounding auxiliary patient data. (See Appendix G.2.3 and G.2.4.)
True negatives ($n=2$): The model correctly ignored confounding auxiliary patient data. (See Appendix G.2.5 and G.2.6.)
False negatives ($n=2$): The model failed to leverage supportive auxiliary patient data. (See Appendix G.2.7 and G.2.8.)
Auxiliary patient data sources—including the indication and history sections, triage data, and reconciled medicines—collectively contributed to the model’s predictions. No single source consistently dominated in providing evidence, with the interplay between these sources frequently complementing one another. A critical challenge for the model lies in its ability to appropriately balance the auxiliary patient data evidence with imaging evidence, particularly when conflicting signals are present. To address this limitation, we propose two key improvements: increasing the size of the training dataset, which is currently relatively small, and adopting an LLM-based decoder. LLMs offer advanced reasoning capabilities, enabling them to better synthesise and prioritise evidence from diverse sources.
@@
Table 3 compares different methods for converting value and category columns into embeddings using the triage and reconciled medicines table, as these contain multiple value and category columns. The aforementioned method of producing embeddings by grouping data from value and category columns (‘Grouped embeddings’) is compared to two other methods. The first is separate embeddings for each datum, where each value column datum is separately transformed using the previously described FNN, while each category column datum is converted to an embedding using a learnable weight matrix, akin to how token embeddings are produced (‘Separate embeddings’). The second method modifies ‘Separate embeddings’ by instead converting the value column data to text and using the decoder’s tokenizer and token embeddings (‘Values-to-text, categories-to-tokens’). The results indicate that the grouped embeddings method was the best representation of heterogeneous patient data for a multimodal language model.
@@ 7 Conclusion
This paper demonstrates the value of incorporating diverse patient data into automated CXR report generation. By integrating patient data from the MIMIC-CXR and MIMIC-IV-ED datasets, we have shown significant improvements in the diagnostic accuracy of generated radiology reports. Our empirical evaluation uncovers new sources of patient information that enhance CXR report generation, including data from ED stays, triaging information, aperiodic vital signs, medicines, and the history section of radiology reports. We present specific methods to convert multimodal patient data into embeddings for a language model, encompassing numerical, categorical, textual, temporal, and image data. We encourage further research and experimentation using our released dataset splits, code, and model checkpoints to explore innovative methods for multimodal patient data integration, with the ultimate goal of enhancing diagnostic accuracy and patient care.
Table 3: Patient data embedding strategies. Underlined indicates a stat. sig. difference to ‘Baseline’ ($p<0.05$).
Embeddings RG CX CB BS
@@
Baseline 29.00 25.81 59.04 23.85
Images + triage + reconciled medicines
Grouped embeddings 31.69 26.72 64.01 24.38
Separate embeddings 25.28 25.32 46.29 23.51
Values-to-text, categories-to-embeddings 30.70 26.46 58.62 24.58